OpenStudy (anonymous):

How do you write a exponent in word form!!!! like if it was x2

OpenStudy (anonymous):

x cubed ?

OpenStudy (baldymcgee6):

Do you mean x squared?

OpenStudy (anonymous):

squared sorry*

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

\[x^2 , \] squared

jimthompson5910 (jim_thompson5910):

To write \(\Large x^2\), you type x^2

OpenStudy (anonymous):

@jim_thompson5910 How'd you change the size of the font ?

jimthompson5910 (jim_thompson5910):

use \large, \Large, or \LARGE

jimthompson5910 (jim_thompson5910):

those are a few font sizes

OpenStudy (anonymous):

Oh.. thanks :D

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

can u help me wth a problem

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

x+6 x-4 ____ x _____ x^2+2x-24 2

jimthompson5910 (jim_thompson5910):

this \[\Large \frac{x+6}{x^2+2x-24}\times\frac{x-4}{2}\] ???

OpenStudy (anonymous):

yess

jimthompson5910 (jim_thompson5910):

ok one sec

OpenStudy (anonymous):

is there a online calculater i can do that on?

jimthompson5910 (jim_thompson5910):

not really, none that I know of

jimthompson5910 (jim_thompson5910):

The idea is to factor everything as much as possible. Then cancel any common factors. \[\Large \frac{x+6}{x^2+2x-24}\times\frac{x-4}{2}\] \[\Large \frac{(x+6)(x-4)}{(x^2+2x-24)(2)}\] \[\Large \frac{(x+6)(x-4)}{2(x^2+2x-24)}\] \[\Large \frac{(x+6)(x-4)}{2(x+6)(x-4)}\] \[\Large \frac{\cancel{(x+6)}(x-4)}{2\cancel{(x+6)}(x-4)}\] \[\Large \frac{(x-4)}{2(x-4)}\] \[\Large \frac{\cancel{(x-4)}}{2\cancel{(x-4)}}\] \[\Large \frac{1}{2}\] ============================================================================ So \[\Large \frac{x+6}{x^2+2x-24}\times\frac{x-4}{2}\] simplifies to \[\Large \frac{1}{2}\] This means that \[\Large \frac{x+6}{x^2+2x-24}\times\frac{x-4}{2}=\frac{1}{2}\] is true for all allowed values of x.

OpenStudy (anonymous):

its so confuusingg:(((( thank u tho!!!!

jimthompson5910 (jim_thompson5910):

where are you lost?

OpenStudy (anonymous):

the whole thing i jus suck at math

jimthompson5910 (jim_thompson5910):

In the first step, I'm just stating the problem

jimthompson5910 (jim_thompson5910):

In the second step, I'm combining the fractions through multiplication

jimthompson5910 (jim_thompson5910):

In step 3, I then rearrange the terms in the denominator

OpenStudy (anonymous):

how do u type the problem like that?

jimthompson5910 (jim_thompson5910):

In step 4, I factor the denominator Then in step 5 (and beyond), I cancel out the common factors til I get \(\Large \frac{1}{2}\) for the answer

jimthompson5910 (jim_thompson5910):

if you just want the plain text version (ie no fancy displayed formulas using LaTex), then [ (x+6)/(x^2+2x-24) ] * [ (x-4)/2 ] [(x+6)(x-4)]/[2(x^2+2x-24)] [(x+6)(x-4)]/[2(x+6)(x-4)] 1/2 ================================================= So [ (x+6)/(x^2+2x-24) ] * [ (x-4)/2 ] simplifies to 1/2

OpenStudy (anonymous):

thank u!!!!:]]]

jimthompson5910 (jim_thompson5910):

yw